Arranged the following compounds according to increasing copper content: Cu2O, CuCO3, Cu2SO4, Cu(MnO4)2

1
Expert's answer
2021-11-09T10:24:56-0500

M(Cu) = 63.54 g/mol

1. M(Cu2O) = 143.09 g/mol

Proportion:

143.09 – 100 %

"63.54 \times 2" – x

"x = \frac{63.54 \times 2 \times 100}{143.09} = 88.81 \; \%"

2. M(CuCO3) = 123.55 g/mol

Proportion:

123.55 – 100 %

63.54 – x

"x = \frac{63.54 \times 100}{123.55} = 51.43 \; \%"

3. M(Cu2SO4) = 223.15 g/mol

Proportion:

223.15 – 100 %

"63.54 \times 2" – x

"x = \frac{63.54 \times 2 \times 100}{223.15} = 56.95 \; \%"

4. M(Cu(MnO4)2) = 301.41 g/mol

Proportion:

301.41 – 100 %

63.54 – x

"x = \frac{63.54 \times 100}{301.41} = 21.08 \; \%"

Answer: Cu(MnO4)2 → CuCO3 → Cu2SO4 → Cu2O